OpenStudy (anonymous):

A merchant has coffee worth $30 a pound that she wishes to mix with 80 pounds of coffee worth $90 a pound to get a mixture that is worth $70 a pound. How many pound of the $30 coffee should be used?

OpenStudy (anonymous):

Let X = no. of pounds of coffee worth $30 The working equation: 30X + 90(80) = 70(80 + X) Solve for X: 7200 - 5600 = 70X - 30X 1600 = 40 X 40 = X Therefore, 40 pounds were used that is worth $30